Circular No. 8256

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

MARS D. C. Parker, Coral Gables, FL, writes that CCD images taken on Dec.

13 UT show that a large regional dust storm has broken out on Mars. A bright dust cloud covers most of the Chryse Planitia and extends west into Candor; most of Aurorae Sinus is obscured. In addition, the dust cloud has extended south into Eos and Margaritifer Sinus. Secondary clouds are seen over Bosporos Planum and northern Argyre and possibly Aram Chaos (on the evening limb). The dimensions of the main cloud are estimated to be ∼ 3000 km east-west and ∼ 1800 km north-south.

V838 MONOCEROTIS M. Mikolajewski and T. Tomov write that the additional star close

to the position of V838 Mon, reported on IAUC 8253, is not a real object. Most probably it is an artifact caused by the extremely red color of the star in combination with a red leak in the U filter, shifted by the differential atmospheric refraction. The two images lie almost along the parallactic angle, and their shift of ∼ 1′′ agrees with the predicted shift (e.g., Filippenko 1982, PASP 94, 715).

COMET C/2002 T7 (LINEAR) Extension to ephemeris on IAUC 8196 (elements MPC 50354):

2003 /04 α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Dec. 7 2^{h}40.^{m}38 +35^{o}06.^{′}6 1.575 2.471 148.^{o}7 12.^{o}0  9.4
17 1 59.59 +31 52.5 1.562 2.335 131.7 18.3  9.2
27 1 25.81 +28 04.0 1.600 2.197 114.4 24.1  9.1
Jan. 6 1 00.17 +24 20.6 1.673 2.056  98.1 28.3  9.0
16 0 41.67 +21 05.4 1.763 1.913  83.1 30.7  8.8
26 0 28.73 +18 25.4 1.854 1.766  69.4 31.4  8.7
Feb. 5 0 19.80 +16 17.9 1.934 1.617  56.6 30.6  8.5
15 0 13.61 +14 36.3 1.991 1.465  44.7 28.3  8.2
25 0 09.15 +13 13.5 2.017 1.311  33.5 24.6  7.9
Mar. 6 0 05.60 +12 01.8 2.004 1.156  22.8 19.5  7.5
16 0 02.30 +10 52.9 1.943 1.002  13.4 13.3  6.9
2003 December 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green